# arXiv:2603.00312v1[cs.AI] 27 Feb 2026

### v1

@@ 1 Introduction
Figure 1 shows our overall approach, which we illustrate with a concrete example of an ECG reasoning trace: “The RR intervals are irregularly irregular, which implies Atrial Fibrillation”. To validate this assertion, we must first verify (i) Perception: Are the R-peaks in the actual signal irregularly spaced? Then, we must verify (ii) Deduction: Do irregular RR intervals necessarily imply Atrial Fibrillation? For Perception, we develop an agentic framework in which an LLM equipped with specialized tools writes code that empirically verifies that the temporal structures described are in fact present in the signal. Next, we evaluate Deduction by comparing the reasoning assertions to the established medical literature. Leveraging the fact that medical diagnoses are exhaustively documented, we first parse online cardiology resources into a structured reference database. Using a Retrieval approach, we can then quantify the validity of a reasoning trace by measuring how often it can retrieve criteria are tagged with the same label as its ground truth label.
@@ 3.1 Perception Phase
That being said, applying this approach is not straightforward due to the deep domain expertise involved in designing data science agents. An innocuous statement, such as “the RR interval are irregularly irregular”, while straightforward to a physician, may not be obvious to those without a background in ECGs. It would require the data science agent to recognize that RR intervals represent the time between R peaks and that R peaks are the highest points of an ECG signal, representing the rapid electrical activation (depolarization) of the heart’s ventricles, in addition to understanding the normal thresholds for regularity.
@@
A) Multi-task Single-Agent. We implement a single-agent system that executes two distinct tasks within a single prompt context: (i) extraction and (ii) verification. First, the agent parses the unstructured reasoning trace into discrete, verifiable findings (e.g. [“inconsistent RR intervals”, “ST elevation >> 0.1mV in V1”]). As shown in Table 2, extraction significantly improves performance by forcing the model to explicitly plan its verification logic before generating code, while simultaneously offering granular interpretability.
@@ Appendix C Knowledge Sources
The aggregation, parsing, and subsequent release of our structured database are strictly permissible under the open-access frameworks governing these sources. As detailed be;pw, LITFL , Wikipedia , and WikiEM operate under CC BY-NC-SA 4.0 or CC BY-SA 4.0, while ECGpedia operates under CC BY-NC 3.0 NL. These Creative Commons licenses explicitly grant permission to ”adapt” (i.e., parse and restructure) and ”share” the underlying content. By releasing our derived database as an open-source asset for non-commercial research with proper attribution, we fully comply with the ShareAlike and NonCommercial provisions, ensuring our work respects the legal terms and open-source ethos of the original contributors. Please see copies of the license in our codebase under the folder license/.
  • LITFL ECG Library: Gold standard for visual ECG recognition and diagnosis, commonly used by cardiologists and electrophysiologists. Licensed under CC BY-NC-SA 4.0, allowing for free adaption and sharing.
  • Wikipedia: Represents the broadest tier of medical consensus with general definitions and widely accepted standard-of-care descriptions. Licensed under CC BY-SA 4.0, allowing for free adaption and sharing.
  • ECGpedia: A specialized knowledge base dedicated designed for teaching with textbook and course materials. Licensed under CC BY-NC 3.0 NL, allowing for free adaption and sharing.
  • WikiEM: Focused on Emergency Medicine, this resource captures rapid triage and decision-making protocols. Licensed under CC BY-SA 4.0, allowing for free adaption and sharing.

### v2

# How Well Do Multimodal Models Reason on ECG Signals?

@@ 1 Introduction
Figure 1 shows our overall approach, which we illustrate with a concrete example of an ECG reasoning trace: “The $R$ intervals are irregularly irregular, which implies Atrial Fibrillation”. To validate this assertion, we must first verify (i) Perception: Are the R-peaks in the actual signal irregularly spaced? Then, we must verify (ii) Deduction: Do irregular $R$ intervals necessarily imply Atrial Fibrillation? For Perception, we develop an agentic framework in which an LLM equipped with specialized tools writes code that empirically verifies that the temporal structures described are in fact present in the signal. Next, we evaluate Deduction by comparing the reasoning assertions to the established medical literature. Leveraging the fact that medical diagnoses are exhaustively documented, we first parse online cardiology resources into a structured reference database. Using a Retrieval approach, we can then quantify the validity of a reasoning trace by measuring how often it can retrieve criteria are tagged with the same label as its ground truth label.
@@ 3.1 Perception Phase
That being said, applying this approach is not straightforward due to the deep domain expertise involved in designing data science agents. An innocuous statement, such as “the $R$ interval are irregularly irregular”, while straightforward to a physician, may not be obvious to those without a background in ECGs. It would require the data science agent to recognize that $R$ intervals represent the time between R peaks and that R peaks are the highest points of an ECG signal, representing the rapid electrical activation (depolarization) of the heart’s ventricles, in addition to understanding the normal thresholds for regularity.
@@
A) Multi-task Single-Agent. We implement a single-agent system that executes two distinct tasks within a single prompt context: (i) extraction and (ii) verification. First, the agent parses the unstructured reasoning trace into discrete, verifiable findings (e.g. [“inconsistent $R$ intervals”, “ST elevation >> 0.1mV in V1”]). As shown in Table 2, extraction significantly improves performance by forcing the model to explicitly plan its verification logic before generating code, while simultaneously offering granular interpretability.
@@ Appendix C Knowledge Sources
The aggregation, parsing, and subsequent release of our structured database are strictly permissible under the open-access frameworks governing these sources. As detailed be;pw, LITFL , Wikipedia , and WikiEM operate under $C$ BY-NC-SA 4.0 or $C$ BY-SA 4.0, while ECGpedia operates under $C$ BY-NC 3.0 NL. These Creative Commons licenses explicitly grant permission to ”adapt” (i.e., parse and restructure) and ”share” the underlying content. By releasing our derived database as an open-source asset for non-commercial research with proper attribution, we fully comply with the ShareAlike and NonCommercial provisions, ensuring our work respects the legal terms and open-source ethos of the original contributors. Please see copies of the license in our codebase under the folder license/.
  • LITFL ECG Library: Gold standard for visual ECG recognition and diagnosis, commonly used by cardiologists and electrophysiologists. Licensed under $C$ BY-NC-SA 4.0, allowing for free adaption and sharing.
  • Wikipedia: Represents the broadest tier of medical consensus with general definitions and widely accepted standard-of-care descriptions. Licensed under $C$ BY-SA 4.0, allowing for free adaption and sharing.
  • ECGpedia: A specialized knowledge base dedicated designed for teaching with textbook and course materials. Licensed under $C$ BY-NC 3.0 NL, allowing for free adaption and sharing.
  • WikiEM: Focused on Emergency Medicine, this resource captures rapid triage and decision-making protocols. Licensed under $C$ BY-SA 4.0, allowing for free adaption and sharing.